OpenStudy (anonymous):

r(t)=(1-t)<-2,-1>+t<1,2> the answer is < -2+3t,-1+3t> but i can't understand how can we do this calculation

OpenStudy (amistre64):

multiply ?

OpenStudy (anonymous):

(-2(1-t), -1(1-t)) +(t, 2t)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ordered pairs.. what i guessed

OpenStudy (anonymous):

=(-2+2t+t,-1+t+2t) =(-2+3t, -1+3t)

OpenStudy (amistre64):

i cant make out what the quaestion is asking for lol

OpenStudy (amistre64):

maybe its the addition of 2 vectors from initial points? or the distance between then or something....

OpenStudy (amistre64):

scalars and addition....ok got it

OpenStudy (anonymous):

it is in line integral notes

OpenStudy (amistre64):

it is vector notation. the () are equations to calculate the scalar functions of each vector; hence(1-t) and (t) cahnge the size of the arrow that represents each vector..

OpenStudy (amistre64):

6<2,3> is an arrow from the origin to a point (2,3); the scalar says to multiply those values by 6 in this case... <6(2),6(3)> = <12,18>which is an arrow from the origin to a point (12,18)

OpenStudy (amistre64):

(1-t)<-2,-1> means to scale the vector by whats in the (): <-2(1-t),-1(1-t)> = <-2+t,-1+t>

OpenStudy (amistre64):

vector t<1,2> is scaled to : <1t,2t>.... now add the vectors together

OpenStudy (amistre64):

<-2+t, -1+t> < 1t , 2t > ------------- <-2t+2t, -1+3t>

OpenStudy (amistre64):

gotta read thru the typos lol

OpenStudy (anonymous):

now u really need some vitamin :P

OpenStudy (anonymous):

-2(1-t) why that's answer isn't -2+2t :S you write -2+t

OpenStudy (anonymous):

oh yes... i mistake, u r right

OpenStudy (amistre64):

you gotta read it for what it is; not for the typos in it lol

OpenStudy (amistre64):

if i type 2+6=9....you gotta have sense enough to understand that im an idiot and look past the fat fingers to the message itself lol

OpenStudy (anonymous):

how sweet ..... suppose u r a teacher, n marking the scripts, underlining the mistakes, n get the same reply from the student ;)

OpenStudy (anonymous):

just kidding :P

OpenStudy (anonymous):

here most of the things must be understood :)

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

i think my question can't replied but thanks

OpenStudy (anonymous):

oh no... u may ask where r u confused

OpenStudy (anonymous):

just go on :)

OpenStudy (amistre64):

If you know about vectors and scalars of vectors and addition of vectors; then the question is answered...

OpenStudy (amistre64):

n<x,y> + m<p,q> <nx ,ny> +<mp,mq> ---------- <nx+mp, ny+mq>

OpenStudy (anonymous):

thanks:D